OpenStudy (anonymous):

What conjecture can you make about the sum of the first 10 positive even numbers? (1 point)The sum is 9 · 10. The sum is 10 · 10. The sum is 11 · 12. The sum is 10 · 11.

OpenStudy (anonymous):

umm hi

OpenStudy (anonymous):

whta is the help needdhere

OpenStudy (anonymous):

i dont know how to find the cojecture of 10 how do you do this

OpenStudy (anonymous):

the problem they gave me to go along with this situation is 2=2=1x2 2+4=6=2x3 2+4+6=12=3x4 2+4+6+8=20=4x5 2+4+6+8+10=30=6x5 how do i find it and those are the answers in the question